In a Δ ABC,if 2 ∠A=3 ∠B=6 ∠C,then the measure of the smallest angle is?

Grade:7

2 Answers

Arun
3 years ago
Let angle C = x
 
Hence angle A = 3x
Angle B = 2x
 
Hence A +B + C = 180°
 
6x = 180°
x = 30° = angle C
Angke B = 60
Angle A = 90°
Vikas TU
3 years ago
Dear student 
2angleA = 3 angle B
A = 3B/2
3B = 6C
C = B/2
Now
A + B + C = 180° ( angle sum property of ∆ )
3B/2 + B + B/2 = 180°
=> (3B+2B+B)/2 = 180°
=> 6B / 2 = 180°
=> 6B = 180×2
=> 6B = 360°
=> B = 360/6
=> B = 60°